# Exploring Young Stellar Variability with Gaia DR3 light curves

### v1

@@ 3.5 Selection of highly variable sources
YSO candidates with GDR3 light-curves were all selected as variable by Gaia pipeline. However, our initial development tests showed that quasiperiodic symmetric variability largely dominates the lowest-amplitude light curves.
Since we wanted to test the applicability of the Q&M indices to a sample with a more significant relative incidence of variability modes, we split our sample in two: a primary sample with only the high-variability sources; and a secondary sample with all other sources passing the previous steps. The primary sample was used to develop the Q&M computation method.
We defined a threshold for selecting the most significant variables as a function of magnitude based on the intrinsic variability of sources in the GAPS sample, which does not necessarily includes only variable sources. A subsample of 479 334 GAPS sources passed all preprocessing steps described in this section. Fig. 3 shows the standard deviation (std) distribution as a function of magnitude for this sample, where std is our proxy of variability for this section. We derived a threshold based on the 97th percentile of the GAPS distribution, estimated by binning the GAPS std distribution using mag bins from 9.1 to 21.9 mag. Each magnitude bin is required to contain at least 500 sources, and bin sizes were sometimes increased by a factor of 1.1 until this minimum number of sources was satisfied. The 97th percentile was then calculated within each bin, and a Savitzky-Golay filter (Savitzky & Golay, 1964) was applied to smooth it, resulting in the percentile curve shown in Fig. 3. This percentile curve was used to select 10,739 highly variable YSOs ( of the initial YSO sample) as part as the primary sample. A secondary sample of 24,054 YSOs below the 97th percentile is also discussed.
@@ 4.1 Asymmetry and Periodicity metrics
We explored the applicability of the periodicity (Q-index) and asymmetry (M-index) variability indices to characterise the GDR3 light curves. The Q-index is defined as (Cody et al., 2014; Hillenbrand et al., 2022):
where is the variance of the residuals of the signal, is the variance of the original light curve, and is the typical uncertainty (calculated here as the average of all squared errors associated with the data). The Q-index is thus a measure of the ratio of the variance of the residuals to the light curve’s general variance. Lower Q values close to zero indicate high periodicity. In contrast, higher Q values (close to unity) reflect irregular or stochastic variability, with no apparent dominant periodic behaviour. The calculation of the Q-index assumes that the light curve contains a signal that may repeat at a characteristic timescale. This signal can be modelled by phase-folding the light curve to this characteristic timescale and employing a smoothing technique to model the signal’s average waveform. A residual curve is then calculated as the difference between the raw-light curve and this waveform.
The M-index is defined as:
where is the mean of the top and bottom deciles of the light curve, is the median of all data points, and is the standard deviation. The M index measures the asymmetry of the light curve by comparing the deviations of the top and bottom deciles relative to the median brightness. Using data points in magnitude, light curves with dipping behaviour have M greater than 0.25, those with bursts have M smaller than (it would be the opposite if the data were in flux), and symmetric light curves will have M-index values around 0.
@@ 4.2 Characteristic timescale
The capital step to quantify the periodicity with the Q-index is to derive a characteristic timescale of variability. Fig. 4 summarises the steps and decision making for timescale derivation, detailed further in this section.
@@ 4.2.1 Generalised Lomb-Scargle periodogram
As a first step, we used Generalised Lomb-Scargle periodograms (Lomb, 1976; Scargle, 1982; Zechmeister & Kürster, 2009)222We employed the Python implementation provided by Astropy (Astropy Collaboration et al., 2018; VanderPlas, 2018) to search for periodic variability in our sample and measure their period. We searched for periods between twice the median time step of the light curve (0.5 days) and the entire light curve duration (900 days). The frequency of the highest peak in the periodogram is selected as the most representative timescale of the light curve variability.
To evaluate the reliability of the timescales obtained with the GLS, we considered a false alarm probability level (FAL) of 1%, computed based on the method by Baluev (2008). Sources without a peak above the FAL were considered aperiodic. Fig. 5 shows examples of GLS periodograms for a periodic and an aperiodic source. A total of 24 067 sources ( of the YSO sample) had their periodicity timescale successfully measured with the GLS. 3 721 of these timescales were longer than half of the total time span of the light curve. These were stored as evidence of long-timescale variability (long-trend) but were not used in the Q-index evaluation.
@@ 4.2.2 Structure Function
In the previous section, 10 726 sources did not show any peak above the FAL in their GLS indicative of periodicity. Although these sources are likely aperiodic variables, their study with the Q-index still required the inference of a characteristic timescale for their aperiodic variability. In their original implementation, Cody et al. (2014) employed the PeakFind (Findeisen et al., 2015) algorithm for aperiodic timescale inference, which requires high cadence sampling to resolve detailed inflexions in light curves (Cody et al., 2022) and thus unattainable with sparsely sampled data sets such as GDR3.
Alternatively, we infer the timescale of aperiodic variability using structure functions (SF). SFs are widely used in Active Galactic Nucleus (AGN) variability studies (Simonetti et al., 1985; Kozłowski, 2016) but have also been applied to variable stars (Eyer & Genton, 1999) and YSOs (albeit in the context of high-cadence surveys; Venuti et al., 2021; Lakeland & Naylor, 2022) to constrain variability timescales. In the context of Gaia DR2, structure functions were successfully used to investigate short timescale variables (Roelens et al., 2017).
Following Hughes et al. (1992) and Roelens et al. (2017), the SF is defined as:
where is the number of pairs separated by a time difference , and the magnitude difference for each pair .
Each bin corresponds to a range of time differences , and is represented by its midpoint . We selected all pairs of observations whose time separation lies within this range to compute SF at each bin.
We discretised the SF in Eq. (3) using logarithmically spaced time difference bins. To further adapt this binning strategy to unevenly sampled light curves, we introduced a tolerance factor , which ensures that each bin contains enough pairs for a reliable calculation (Roelens et al., 2017). Bins that do not contain any pairs are excluded from the analysis. If a populated bin initially has too few pairs, its size is increased symmetrically until the minimum requirement of at least three pairs is met. With GDR3 data, we could compute SFs for time differences in the range 0.07–980 d (see Fig. 31). The number of bins used was based on the number of epochs available for each light curve, with 1.5 times the number of epochs used for the number of bins.
For aperiodic variability, the SF typically shows three main regimes. A noise-dominated regime, where the SF remains roughly constant around the intrinsic noise level in the data. A variability regime, where the SF value increases until it peaks at the turnover timescale . And a plateau, where beyond , the SF flattens as no new variability occurs. To represent these three regimes, we fitted the SFs with a function of the form:
@@
We fit Eq. (4) to the SF using the MIGRAD
routine (James & Roos, 1975) implemented in Python as part of the iminuit
library (Dembinski & et al., 2020). Initial guesses and limit values for the model parameters were based on inspection of the distribution of all SF shown in Fig. 31 and were chosen as follows: , , d for the initial guess.
While is in the range mag2, in mag2 and in d. Examples of SF model fit results can be seen in Fig. 6.
The imposition of model parameter limits resulted in a relative excess of fit results with edge values. Hence, we filtered these results to remove spurious models within 5% from the edge values for , , and for the lower limit of . A total of 5 354 SF model fits were discarded at this step. Visual inspection of discarded SF models showed that these were typically sources with complex multimode variability, including different timescales and that cannot be successfully modelled by Equation (4), or sources with poorly sampled SF with a long time range without data (blind time range, further discussed below). An excess of upper limit values of also existed, and we verified that these were typically related to light curves exhibiting a long-term trend (e.g., due to a gradual change in brightness) with a timescale longer than the timespan of GDR3 observations, where the SF continues to increase without ever reaching the plateau, in which case, we stored the timescale as evidence for long-term variability, but did not use them for Q-index evaluation.
In addition to the issue of edge values, 12 best-fit models were also discarded because they had limited sampling around the derived timescale. This issue is further discussed in the App. B, and it is introduced by GDR3 sparse time-series sampling, which for some sources results in a blind time range in the SF. For affected sources (see Fig. 6-bottom), no data points exist in the range days. Beyond this range, the SF show a flat distribution - hence, it is impossible for models to correctly identify the transition from the variability regime into the plateau.
The fits of the SF model in this section allowed us to estimate the aperiodic variability timescale for sources. For sources, the SF resulted in timescales in upper-edge values for , or in timescales that were longer than half of the total time span of the light curve, and these were stored as evidence of long-trend but were not used in the Q-index evaluation. For , we were unable to derive a time scale of variability using the GLS or SF methods.
@@ 4.3 Residual spurious variability
Following the discussion in App. 29, we further remove 1 251 sources from our sample as their measured timescale coincided with typical timescales introduced by the spurious variability introduced by Gaia’s scan-angle effects (Sect. 3.3). After this final cleaning step, our final sample YSO light curves further analysed were composed of 28 179 sources, with 9 824 of these in the primary sample, and 18 355 in the secondary sample. We only analysed sources for which we could derive a timescale.
@@ 4.4 Q and M implementation and morphology classes
For the Q-index, once a characteristic timescale is derived, the light curve is folded to its corresponding phase. Next, we derive its waveform using a Savitzky-Golay filter with a window size of 25% of the phase and a second-order polynomial. This set of parameters allowed us to obtain waveforms representative of the ongoing variability while avoiding overfitting. The estimated waveform is then extracted from the data to obtain the residual curve. Finally, Q-indices can then be estimated with Eq. (1).
Following the seminal study by Cody et al. (2014), the Q and M indices can be used to gauge light-curve morphologies and sort variables into seven standard classes:
Eclipsing binary (EB): and ;
Strictly Periodic (P): and ;
Quasi-periodic symmetric (QPS): and ;
Quasi-periodic dippers (QPD): and ;
Aperiodic dippers (APD): and ;
Bursters (B): and ;
Stochastic (S): and ;
Previous literature included additional classes such as multiperiodic (MP), non-variable (NV) and long-trend (L). These classes are not based directly on the Q and M indices, but are defined based on the failure to apply Q&M methodology. The identification of multiperiodic variability in previous studies relied on the high cadence of their observations and required additional steps, including visual inspection of the light curves. We did not reproduce those steps as part of this study. The NV class was also excluded, since our sample is focused on light curves with enhanced variability. Long-trend variables were identified here as sources with inferred timescale longer than half of GDR3 timespan.
@@ 5 Results
5.1 Asymmetry and Periodicity indexes
@@
Following the filtering steps in Sect. 3 and the methods in Sect. 4 we successfully derived Q-indices for 23 213 sources (8 385 in the primary sample) and M-indices for 28 179 sources (9 824; primary). Examples of light curves at different Q&M values are shown in Fig. 7 and Fig. 8, illustrating how we successfully adapted the Q&M methodology for use with Gaia DR3 light curves. In Fig. 7, light curves at low Q values are very well represented by the estimated periodic waveform. At intermediate Q-values, periodic waveforms are still found, but with an increased phase dispersion. At Q values close to unity, light curves are dominated by stochastic variability, where the timescales inferred can no longer be interpreted as a periodic timescale and the data points are stochastically distributed around the average. In Fig. 8, light curves with extreme asymmetries have M values farthest away from 0, with negative M values corresponding to variables with brightening events and positive M values corresponding to dimming events, while more symmetric light curves are closest to .
@@
Fig. 9 shows the distribution of Q versus M for our sample. This general distribution peaks at 0.43 (quasi-periodic) and 0.05 (symmetric). The individual distributions of the Q and M indices are shown in Fig. 10, where we separate our sample by timescale method. As expected from the nature of the methodology involved in the derivation of variability timescales, the distribution of Q for sources with a periodic timescale derived by the GLS (in red) is mainly on the periodic/quasi-periodic region of the diagram, peaking at , with the Q-index distribution for sources with aperiodic timescales derived from SFs peaking at . For the M-index, sources with periodic timescales peak at 0.02, and at 0.15 for aperiodic timescales, with the aperiodic distribution showing a tail towards dipping M-index values. Fig. 10 also compares the distribution of Q&M-indices for our primary and secondary samples, where the differences for these two samples are also supported by Kolmogorov–Smirnov tests. Quasi-periodic variability is the most frequent mode among our sample, accounting for more than 69% of the sample. The secondary sample peaks at and , being dominated by quasi-periodic symmetric sources - as mentioned in Sect. 3.5 the over-weight of this population of low-amplitude quasiperiodic symmetric sources on the general occurrence of variability classes was the motivation to discuss results as a function of a higher amplitude vs lower amplitude variable samples. The primary sample of variables peaks at and , with a more relevant proportion of dipping and aperiodic variables in comparison to the secondary sample.
The classification scheme outlined in Sect. 4.4 was applied to classify light curves into seven morphological classes with an eighth class to account for long-trend variables. The count and frequency of sources in each class are summarised in Table 3. As already suggested by the Q&M distributions, dipper variability is the most common among the primary sample, with 38% of the primary sample showing aperiodic (APD) or quasi-periodic (QPD) dipper variability.
@@
We also identified a total of 4 966 long-trend (L) variables, where the characteristic timescale derived was longer than half the total time span of the GDR3 observations (typically days). These account for 3 721 periodic timescales derived with the GLS and 1 45 aperiodic timescales derived with the SF. We visually inspected a large fraction of these light curves and confirmed that a long-term trend was indeed present. These timescales are at the limit of GDR3 resolvable timescales and should be interpreted as a bottom limit for the real variability timescale, and hence we did not employ them for Q-index calculation, although we discuss them as part of our results.
Full sample | Primary | Secondary | |||||||||||||
Frequency | ts- | Frequency | ts- | Frequency | ts- | ||||||||||
Class | count | Disk | CTTS | ratio | ptp | count | Disk | CTTS | ratio | ptp | count | Disk | CTTS | ratio | ptp |
P | 286 | 17.1 | 13.7 | 13.4 | 0.19 | 199 | 17.6 | 14.6 | 26.6 | 0.22 | 87 | 11.5 | 8.7 | 5.8 | 0.10 |
QPS | 9 514 | 20.2 | 14.8 | 2.6 | 0.15 | 2 362 | 42.1 | 36.6 | 1.4 | 0.31 | 7 15 | 13.0 | 6.1 | 3.3 | 0.09 |
QPD | 5 439 | 39.4 | 24.7 | 1.8 | 0.35 | 2 886 | 62.1 | 40.6 | 1.3 | 0.57 | 2 553 | 13.7 | 4.5 | 2.5 | 0.09 |
APD | 1 103 | 56.9 | 39.7 | 0.4 | 0.53 | 822 | 68.4 | 48.2 | 0.4 | 0.67 | 281 | 23.5 | 3.0 | 0.7 | 0.13 |
B | 5 252 | 21.7 | 22.0 | 2.0 | 0.16 | 1505 | 45.2 | 45.6 | 1.3 | 0.31 | 3 747 | 12.2 | 7.0 | 2.4 | 0.10 |
S | 1 490 | 28.2 | 23.6 | 0.8 | 0.19 | 601 | 43.3 | 49.0 | 0.4 | 0.30 | 889 | 18.0 | 7.4 | 1.1 | 0.12 |
EB | 129 | 17.1 | 11.4 | 23.6 | 0.20 | 95 | 16.8 | 14.1 | 91.0 | 0.24 | 34 | 17.7 | 0.0 | 6.8 | 0.10 |
L | 4 966 | 21.1 | 11.0 | - | 0.16 | 1 354 | 40.0 | 33.6 | - | 0.37 | 3 612 | 14.1 | 2.5 | - | 0.09 |

### v2

@@ 3.5 Selection of highly variable sources
Initial tests revealed that low-amplitude QPS was the most common variability mode in the YSO sample, dominating the light curves of 41% of the sources analysed (see Sect. 6.2.2 and Table 1). However, to assess the applicability of the Q-M indices to GDR3 light curves, we concentrated our pipeline development on a sample selected to include a higher proportion of less common variability modes. This was achieved by focussing our initial analysis on a sub-sample of the most variable sources.
To support the identification of the most variable sources as a function of magnitude, it was relevant to first understand the behaviour of the YSO sample compared to non-variable stars also observed by GDR3. Although GDR3 does not provide light curves for non-variable YSOs, this comparison was still possible based on the GAPS sample, which includes GDR3 light curves for all sources in the Andromeda field, irrespective of their variability status. The GAPS and YSO samples share GDR3 photometric performance, albeit having different sky locations and numbers of transits (see Fig. 1). A comparison between the two samples as a function of their average G magnitude and standard deviation (std) is provided in Fig. 3, where we display properties of light curves that passed pre-processing steps described in this section (34 793 YSO and 479 334 GAPS light curves), i.e. already accounting for the minimum number of epochs required for stable variability indices (Sect. 3.2, App. A.1). The GAPS sample is predominantly made up of non-variables, with roughly 1% of its sources selected by GDR3 variability processing. This is evident in Fig. 3, where the bulk of the GAPS distribution (yellow regions, top panel) is located below the lower envelope of the YSO sample distribution (derived as the 1st rolling percentile of that distribution).
We defined a threshold for selecting the most significant variables as a function of magnitude based on rolling percentiles of the GAPS distribution, which were estimated by binning the GAPS distribution using mag bins from 9.1 to 21.9 mag. Each magnitude bin was required to contain at least 500 sources, and bin sizes were sometimes increased by a factor of 1.1 until this minimum number of sources was satisfied. Rolling percentiles were then calculated within each bin, and a Savitzky-Golay filter (SavitzkiGolay1964A) was applied to smooth it. The 97th-percentile threshold provided the best compromise between selecting highly variable sources while still retaining a large enough sample of variables.
Finally, we selected 10 739 sources in the YSO sample with a std larger than 97% of GAPS sources in their magnitude bin. These sources are the primary sample on which we focused the adaptation of the Q-M method (Sect. 4). The remaining 24 054 sources below the 97th percentile of the GAPS distribution are hereafter named secondary sample; these were also analysed once the workflow for the Q-M analysis was settled.
@@ 4.1 Asymmetry and periodicity metrics
We explored the applicability of the Q-M indices to characterise the GDR3 light curves. The Q index is defined as (Cody_14; Hillenbrand2022):
where is the variance of the residuals of the signal, estimated by subtracting a waveform estimation (derived by smoothing the phase-folded light curve, folded to the characteristic variability timescale) from the raw light curve. is the variance of the original light curve, and is the square of the typical uncertainty (calculated here as the average of all squared errors associated with the data). The Q index is thus a measure of the ratio of the variance of the residuals to the general variance of the light curve. Lower Q values close to zero indicate high periodicity. In contrast, higher Q values (close to unity) reflect irregular or stochastic variability. The derivation of the Q index assumes that the light curve contains a signal that may repeat at a characteristic timescale. This signal can be modelled by phase-folding the light curve to its characteristic timescale and employing a smoothing technique to model the average waveform of the signal. A residual curve is then calculated as the difference between the raw-light curve and this waveform.
The M index measures the asymmetry of the light curve by comparing how the upper and lower brightness deciles deviate from the median. It is is defined as
where is the mean of the highest and lowest 10% of light-curve points, is the median magnitude, and is the standard deviation. When working in magnitudes, light curves with dipping behaviour have , those with bursts have , and symmetric light curves have .
@@ 4.2 Characteristic timescale
The capital step to quantify the periodicity with the Q index is to derive a characteristic timescale of variability. Fig. 4 summarises our workflow for timescale derivation: We first applied the generalised Lomb-Scargle periodogram to identify periodic timescales, assessing their reliability using the false alarm probability (Sect. 4.2.1). If no reliable periodic timescale was found, we then computed their structure function to derive an aperiodic timescale (Sect 4.2.2). Sources were discarded if neither method provided a trustworthy timescale. A final step removed sources with timescales affected by the typical spurious variability of Gaia (Sect 4.3)
@@ 4.2.1 Generalised Lomb-Scargle periodogram
We used the generalised Lomb-Scargle periodogram (GLS; Lomb1976; Scargle1982; Zechmeister2009; Vanderplas2018) to search for periodic variability in our sample. We searched for periods between twice the median time step of the light curve (0.5 days) and the entire light curve duration (900 days), with the frequency of the highest peak recorded as the most representative timescale of the light curve variability. The reliability of GLS timescales was evaluated considering a false-alarm probability level (FAL) of 1%, with a Baluev2008MNRAS.385.1279B FAL estimation. Sources without a peak above the FAL were considered aperiodic. Fig. 5 shows examples of GLS periodograms for a periodic and an aperiodic source. A total of 24 067 sources ( of the YSO sample) had periodic timescales measured with the GLS. 3 721 of these timescales were longer than half the total time span of the light curve, and were stored as evidence of L type variability, but were not used to derive the Q index.
@@ 4.2.2 Structure function
In the previous section, 10 726 sources did not show any peak above the FAL in their GLS, indicative of periodicity. These sources are likely aperiodic variables. Still, their study with the Q metric required the inference of a characteristic timescale for their aperiodic variability. In their original implementation, Cody_14 employed the PeakFind (Findeisen_2015ApJ...798...89F) algorithm for aperiodic timescale inference, which requires high-cadence sampling to resolve detailed inflexions in light curves (Cody_2022) and thus is unattainable with sparsely sampled data sets such as GDR3. Alternatively, we inferred aperiodic timescales using structure functions (SF). SFs are widely used in other fields to constrain variability timescales, with varied applications in the study of Active Galactic Nucleus (Simonetti85; Kozlowski2016, AGN), other variable stars (1999A&AS..136..421E), and with a few applications to YSOs albeit in the context of high-cadence surveys(Venuti_2021; Lakeland2022). In the context of Gaia DR2, structure functions were successfully used to investigate short timescale variables (Roelens2017).
Following Hughes1992 and Roelens2017, the SF of a light curve is defined as
where is the number of pairs separated by a time difference , and the magnitude difference for each pair . We discretised Eq. (3) using logarithmically spaced bins. Each bin corresponds to a range of time differences , and is represented by its midpoint . We selected all pairs of observations whose time separation lies within this range to compute SF at each bin. Bins that do not contain any pairs are excluded from the analysis. To further adapt this binning strategy to unevenly sampled light curves, we introduced a tolerance factor , which ensures that each bin contains enough pairs for a reliable calculation (Roelens2017). If a populated bin initially has too few pairs, its size is increased symmetrically until the minimum requirement of at least three pairs is met. With GDR3 data, we could compute SFs for time differences in the range 0.07–980 d (see Fig. 33). The number of bins used was based on the number of epochs available for each light curve, with 1.5 times the number of epochs used for the number of bins.
For aperiodic variability, the SF typically shows three main regimes. A noise-dominated regime where the SF remains roughly constant around the intrinsic noise level in the data. A variability regime where the SF value increases until it peaks at the turnover timescale . And a plateau where, beyond , the SF flattens as no new variability occurs. To represent these three regimes, we fitted the SFs with a function of the form
@@
We fit Eq. (4) to the SF using the MIGRAD routine (James:1975dr) implemented in Python as part of the iminuit library (iminuit). The initial guesses and limit values for the model parameters were based on inspection of the distribution of all SF shown in Fig. 33 and were chosen as follows: , , d for the initial guess. While is in the range mag2, in mag2 and in d. Examples of SF model fit results can be seen in Fig. 6. The imposition of model parameter limits resulted in a relative excess of fit results with edge values. Hence, we filtered these results to remove spurious models within 5% of the edge values for , , and for the lower limit of . A total of 5 354 SF fits were discarded at this step. Visual inspection of discarded models showed that these were typically sources with complex multi-mode variability, including different timescales and that could not be represented by Equation (4), or sources with poorly sampled SF with a long time range without data (blind time range). An excess of upper limit values of also existed, and we verified that these were typically related to light curves exhibiting a long-term trend (e.g. a gradual change in brightness) with a timescale longer than the time span of GDR3 observations, where the SF continues to increase without ever reaching the plateau, in which case, we stored the timescale as evidence for long-term variability but did not use them to evaluate the Q index. In addition to the issue of edge values, 12 best-fit models were also discarded because they had limited sampling around the derived timescale. This issue is further discussed in the App. B, and it is introduced by GDR3 sparse time-series sampling, which for some sources results in a blind time range in the SF. For affected sources (see Fig. 6-bottom), no data points exist in the range days. Beyond this range, the SF show a flat distribution - hence, it is impossible for models to correctly identify the transition from the variability regime into the plateau.
The fits of the SF model in this section allowed us to estimate the aperiodic variability timescale for sources. For sources, the SF-fit resulted in upper-edge values or timescales that were longer than half of the total time span of the light curve, with these stored as evidence of long trend, but were not used to evaluate the Q index. For , we were unable to derive a time scale of variability using the GLS or SF methods.
@@ 4.3 Residual spurious variability
We further discarded 1 251 sources as their measured timescale coincided with typical timescales caused by the spurious variability introduced by the Gaia scan-angle effects (Sect. 3.3, see also App. 31). After this final cleaning step, we proceeded with the analysis of 28 179 YSO light curves, with 9 824 of these in the primary sample and 18 355 in the secondary sample.
@@ 4.4 Q and M implementation and morphology classes
Once a characteristic timescale was derived, light curves were folded to their corresponding phase, and a waveform was derived using a Savitzky-Golay filter with a window size of 25% of the phase and a second-order polynomial. This set of parameters allowed us to obtain waveforms representative of the ongoing variability while avoiding overfitting. The estimated waveform was then subtracted from the data to obtain the residual curve, followed by estimating the Q indices with Eq. (1).
Following Cody_14 scheme, the Q and M indices can be used to gauge light curve morphologies and sort variables into seven standard classes:
EB: and .
P: and .
QPS: and .
QPD: and .
APD: and .
B: and .
S: and .
Additional classes are sometimes included in the literature, such as multi-periodic, non-variable, and L. These classes are not based directly on the Q and M indices, but are defined based on the failure to apply Q-M method. The identification of multi-periodic variability in previous studies relies on the high cadence of their observations and requires additional steps, including visual inspection of the light curves. We did not reproduce those steps as part of this study. As our sample is focused on light curves with enhanced variability, the non-variable YSOs are also excluded. L variables were identified here as sources with inferred timescales longer than half of GDR3 time span.
@@ 5 Results
5.1 Asymmetry and periodicity indices
@@
We successfully derived Q indices for 23 213 sources (8 385 in the primary sample) and M indices for 28 179 sources (9 824; primary). Our successful adaptation of the Q-M method for use with GDR3 light curves is illustrated in Fig. 8 and Fig. 9. In Fig. 8, light curves at low Q values are very well represented by the estimated periodic waveform. At intermediate Q values, quasi-periodic waveforms are still found, but with increasing phase dispersion. At , light curves are dominated by stochastic variability, timescales inferred can no longer be interpreted as a periodic timescales, with data points are stochastically distributed around the average. In Fig. 9, light curves with extreme asymmetries have M values farthest away from 0, with negative M values corresponding to variables with brightening events and positive M values corresponding to dimming events, while more symmetric light curves are closest to .
@@
Fig. 7 shows the distribution of Q versus M for our sample. This general distribution peaks at (quasi-periodic) and (symmetric). The individual distributions of the Q and M indices are shown in Fig. 10, where we separate our sample by timescale method. As expected from the nature of the methodology for measuring variability timescales, the distribution of Q for sources with a periodic timescale derived with the GLS (in red) is mainly on the periodic and quasi-periodic regions of the diagram, peaking at , with the Q index distribution for sources with aperiodic timescales derived from SFs peaking at . For the M index, sources with periodic timescales peak at 0.02, and at 0.15 for aperiodic timescales, with the aperiodic distribution showing a tail towards dipping M index values. Fig. 10 also compares the distribution of Q-M indices for our primary and secondary samples, where the differences for these two are also supported by Kolmogorov–Smirnov (KS) tests. Quasi-periodic variability is the most frequent mode in our sample, accounting for more than 69% of sources. The secondary sample peaks at and , being dominated by QPS sources. The over-representation of low-amplitude QPS sources in the general occurrence of variability classes motivated discussing results as a function of higher-amplitude (primary) vs lower-amplitude (secondary) variable samples. The primary sample of variables peaks at and , with a larger proportion of dipping and aperiodic variables in comparison to the secondary sample.
@@
The classification scheme outlined in Sect. 4.4 was applied to sort light curves into seven morphological classes with an eighth class attributed to account for L variables, which had characteristic timescales longer than half the total time span of the GDR3 observations (typically days). The count and frequency of sources in each class are summarised in Table 1. As already suggested by the Q-M distributions, dipper variability is the most common among the primary sample, with 38% of these sources showing APD or QPD variability. We identified 4 966 L variables, 3 721 with periodic timescales derived with the GLS and 1 275 aperiodic timescales derived with the SF. We visually inspected a large fraction of these light curves and confirmed that a long-term trend was indeed present. Their timescales are at the limit of GDR3 resolvable timescales and should be interpreted as a bottom limit for the real timescale, and hence we did not employ them to calculate the Q index, although we discuss them as part of our results.